An 85kg person is on the moon (moon acceleration due to gravity is 1.6ms) Contrast the mass and weight of this person on the moon
1
Expert's answer
2012-10-24T10:59:46-0400
The weight of the person on the moon is

P = m*g = 85[kg]*1.6[m/s²] = 136[N].